Opened 11 years ago

Last modified 2 years ago

#341 assigned flight processing

EU10/01, 168/2010, Loobos — at Version 41

Reported by: jb362 Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors: knpa

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EU10_01-2010_168_loobos

Data arrived from ARSF via network transfer on 23/06/2010

Scientific objective:

1.Key impact factors on CO2 and H2O fluxes
1.1 examine the influence of within-site elevation changes using flux tower observations, airborne LiDAR data, and ecosystem models data (Loobos)
1.2 develop an approach to allocate CO2 and H2O fluxes to specific tree species of a mixed forest (Norunda)
1.3 assess the influence of forest thinning on CO2 uptake (Norunda)
1.4 evaluate above findings and test transferability of findings (Tharandt)
1.5 investigate the impact of storm damage (Tharandt)

  1. Reducing uncertainty in the upscaling of greenhouse gases

2.1. develop an approach to evaluate airborne flux measurements using flux tower data, footprint model and airborne LiDAR data (Loobos)
2.2 use airborne LiDAR data to validate GLAS/ICESat for different forest biomes, where accurate tree height and vegetation cover density will be co-located with GLAS footprints
2.3 evaluate the ability of GLAS spatial sampling density to represent the wider field site
2.4 develop an approach to model CO2 and H2O fluxes on a regional scale using GLAS/ICESat data and generate key driving parameters for land-surface models
2.5 evaluate whether regional CO2/H2O flux models driven by satellite optical data can be improved by the addition of tree height information from satellite LiDAR

Priority: a5

PI: N. Kljun

Sensors:

  • Eagle (01/09/2010)
  • Hawk
  • Leica LIDAR (06/09/10)
  • RCD (26/08/10)

Change History (41)

comment:1 Changed 11 years ago by jb362

This directory needs to be backed up in the repository

comment:2 Changed 11 years ago by jb362

IPAS navigation for this project has been processed.

comment:3 Changed 11 years ago by benj

RINEX data downloaded from Thelma.

comment:4 Changed 11 years ago by ella

Applanix navigation processing now done. Files located in ARSF workspace.

comment:5 Changed 11 years ago by jb362

Started hyperspectral processing

comment:6 Changed 11 years ago by jb362

I need to check the Sbet file times against the flight line times in PosGNSS for this project as they don't seem to match the logsheet, however the extract directory on workspace is empty so there is no saved PosPac/PosGNSS data and no raw files. I think Loobos is going to have to be re-downloaded so that I can put the raw files back into workspace and then open up the project in PosGNSS

comment:7 Changed 11 years ago by jb362

downloading raw files done

comment:8 Changed 11 years ago by jb362

  • Owner set to jb362
  • Status changed from new to assigned

comment:9 Changed 11 years ago by jb362

The -1.nav didn't contain the navigation data but we have used the datacube.nav file and renamed it as -1.nav. Original -1.nav file renamed as -1.nav_old.

comment:10 Changed 11 years ago by jb362

.raw file missing for first eagle line

comment:11 Changed 11 years ago by jb362

No eagle data for flight line 1 or 9 and no hawk data for flight line 9 because of missing .raw files

comment:12 Changed 11 years ago by jb362

Actually .raw files do exits for flight line 9 on each sensor... I don't know why neither produced .tif files

comment:13 Changed 11 years ago by jb362

Offset times:

h1: 0.01
h2: 0.03
h3: 0.01
h4: 0.01
h5: 0.01
h6: 0.03
h7: 0.01
h8: 0.01
h9: Missing flightline
h10: 0.01
h11: 0.01
h12: 0.00
h13: 0.03
h14: -0.01
h15: 0.01
h16: 0.01
h17: 0.01
h18: 0.01
h19: 0.01
h20: 0.02
h21: 0.02
h22: 0.01
h23: 0.01
h24: 0.01
h25: -0.01
h26: 0.01
h27: 0.01
h28: 0.00
h29: 0.01

e1: Missing flightline
e2: 0.03
e3: 0.04
e4: 0.01
e5: 0.01
e6: 0.04
e7: 0.03
e8: 0.05
e9: Missing flightline
e10: 0.01
e11: 0.01
e12: 0.01
e13: -0.01
e14: -0.01
e15: 0.03
e16: 0.02
e17: 0.00
e18: 0.04
e19: 0.03
e20: 0.03
e21: 0.03
e22: 0.00
e23: 0.00
e24: 0.01
e25: 0.04
e26: -0.05
e27: 0.05
e28: 0.01
e29: 0.04

comment:14 Changed 11 years ago by jb362

Hyperspectral complete and waiting for delivery check

comment:15 Changed 11 years ago by chrfi

handing lidar over to knpa

roll offsets found so far

1202250.00093198
1211550.00090198
1342270.00093198
1347580.00093198
1337000.00096198
1326130.00093198
1221440.00093198
1351380.00093198
1358570.00093198
1404410.00093198
1410260.00093198
1416060.00093198
1421410.00090198
1433040.00093198

comment:16 Changed 11 years ago by knpa

  • Other processors set to knpa

Taking over responsibility for LIDAR.

comment:17 Changed 11 years ago by iopa

Performing delivery check on hyperspectral.

comment:18 Changed 11 years ago by iopa

Results:

  1. PI name spelling in Readme front page --> should be N. Kljun
  1. Flight line 9 for eagle and hawk has raw files and lev1's, so should produce .tif's as well.
  1. Since line 1 for eagle does not have a raw file, the rest of the lines will have to shift up so there is a continuous set of lines 1-28 for eagle. For the readme, this means that the lev1's and tif's 1-28 will need to be matched with the recorded line numbers in the logsheet for lines 2-29 (1 will be 004, 2 will be 019 etc.). This therefore applies to lev1's, lev3's and readme which will need to be renamed. (all this assuming eagle line 1 can not have a lev1 and lev3)
  1. Remove any hidden files in the delivery, currently there is just .Read_Me-20100722.tex.swp in EU10-01.
  1. Remove the .tex after the pdf is produced. Also remove the .tex in the screenshots directory.
  1. Missing bin directory, with instructions on downloading azgcorr (see http://arsf-dan.nerc.ac.uk/trac/wiki/Processing/FilenameConventions#Delivereddata)

comment:19 Changed 11 years ago by jb362

Made delivery check fixes

comment:20 Changed 11 years ago by jb362

Additional SCT offsets:

e9a: 0.01
e9b: 0.01
h9a: 1.30
h9b: 1.55

comment:21 Changed 11 years ago by jb362

  • Description modified (diff)

comment:22 Changed 11 years ago by jb362

If hawk data has not been rectified in two weeks time then eagle data should be delivered separately.

comment:23 Changed 11 years ago by jb362

Eagle SCT offsets:
e1: 0.04
e2: -0.01
e3: 0.08
e4: -0.10
e5: -0.05
e6: -0.08
e7: 0.09
e8: -0.18
e9: 0.28
e10: -0.30
e11: 0.09
e12: 0.05
e13: -0.06
e14: -0.10
e15: 1.26
e16: 0.00
e17: 0.03
e18: -0.07
e19: 0.03

The large SCT offset for flight line 15 is due to a sync error.

The header file for eagle flight line 16 had to be corrected so that "lines = 9596", rather than "lines = 1837" as in the original header file.

comment:24 Changed 11 years ago by jb362

Sorry, these offsets are for Santander (!!!)

comment:25 Changed 11 years ago by emca

Starting camera processing

comment:26 Changed 11 years ago by knpa

Line Roll
115710 0.00093198
120225 0.00093198
120723 0.00093198
121155 0.00090198
121650 0.00093198
122144 0.00093198
122708 0.00093198
123230 0.00093198
123725 0.00093198
124233 0.00093198
124752 0.00093198
125357 0.00093198
125940 0.00093198
130510 0.00093198
131013 0.00093198
131607 0.00093198
132108 0.00093198
132613 0.00093198
133147 0.00093198
133700 0.00096198
134227 0.00093198
134758 0.00093198
135318 0.00093198
135857 0.00093198
140441 0.00093198
141026 0.00093198
141606 0.00093198
142141 0.00090198
143304 0.00093198

comment:27 Changed 11 years ago by emca

Camera processing complete.
Removed photographs 1-11 as they did not cover the study area. Also removed 22, 44, 135, 136, 156, 202, 203, 221, 238, 239, 284, 285 & 292 due to under/over exposure.

Ready for delivery check at:
~airborne/workspace/EU10_01-2010_168_Loobos/delivery/20100824

comment:28 Changed 11 years ago by jb362

Delivery checking camera.

comment:29 Changed 11 years ago by jb362

Everything looks fine. Should PI on Readme just be N. Kljun or Dr N. Kljun?

comment:30 Changed 11 years ago by benj

Doesn't really matter. Pick one.

comment:31 Changed 11 years ago by emca

Photography data has been sent to PI 26/08/2010 along with LiDAR and Eagle data from day 168 (ticket #345)

comment:32 Changed 10 years ago by knpa

Finished processing LIDAR, creating delivery.

comment:33 Changed 10 years ago by jb362

Eagle data delivered 01/09/2010.

comment:34 Changed 10 years ago by ella

Starting LiDAR delivery check.

comment:35 Changed 10 years ago by ella

LiDAR check complete.

-Flight logsheet misspells site name.

-Read_me document figure 1 is intensity image without vectors but the caption says it has vectors overlaid.

comment:36 Changed 10 years ago by knpa

Fixed above issues, preparing delivery to PI.

comment:37 Changed 10 years ago by knpa

LIDAR delivered today (06/09/10) to N. Kljun at Swansea University.

comment:38 Changed 10 years ago by knpa

  • Description modified (diff)

comment:39 Changed 10 years ago by knpa

  • Description modified (diff)

comment:40 Changed 10 years ago by jb362

  • Description modified (diff)

comment:41 Changed 10 years ago by knpa

  • Description modified (diff)
  • Owner changed from jb362 to knpa

Has been transferred to repository and can be r-synced.

Note: See TracTickets for help on using tickets.